OpenStudy (anonymous):

A contractor decided to build a bridge in 35 days. He employed 60 men in the beginning and 40 more after 17 days and completed the work on time. If he had not employed the additional men, how many days behind schedule would it have been finished?

OpenStudy (anonymous):

A person divided his total amount into two parts in the ratio of 3:2. He invested the first part in a bank which gives simple interest at the rate of 10% pa and the second part in another bank which gives simple interest at the rate of 20%. If after 3 years the total interest received by him from both the banks is 42% of the total amount he deposited in both the banks together, what is the total amount he deposited?

OpenStudy (kropot72):

We need to find how many man-days were needed to build the bridge. Initially 60 men worked for 17 days = 60 * 17 man-days Then 100 men worked for 18 days = 100 * 18 man-days The total man-days to build the bridge = (60 * 17) + (100 * 18) = you can calculate. When you have posted the total man-days I will explain the final step.

OpenStudy (anonymous):

the answer is total 47 days .so 12 days more.but i didnt understand that 60 men do a work for 18 days remaing 17 days with 100 man?.plse expalin me

OpenStudy (kropot72):

The time to build the bridge = (60 * 17) + (100 * 18) = 2820 man-days. If the original number of men (60 men) was not increased it would take\[\frac{2820}{60}=47\ days\] The contractor employed 60 men for the first 17 days. Then he employed an additional 40 men from day 18 until day 35. Therefore 60 men worked for 17 days and the 60 + 40 = 100 men worked for 18 days.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank u

OpenStudy (kropot72):

You're welcome :)